a carnot engine whose efficiency is 50 %,if sink temperature is decreased by 100% then efficiency will be 70% ,find the temperature of source

Grade:12th pass

1 Answers

Shailendra Kumar Sharma
6 years ago
1-Tsink/TSource =.5
How can you reduce temp by 100% 
still lets say you reduced to half 
1-(.5Tsink)/Tsource =.9
This gives 
(TSource-Tsink)/(Tso-.5Tsi) =5/9
4Tsource=(6.5) Tsink
This way you can proceed